# bibliography.bib

### v1

@@ 3 Local Axisymmetric Black Holes
Chandrasekhar obtained all vacuum spherical topology axisymmetric black holes in Schwarzschild-like coordinates [chandrasekharbook, LocalToroidalXanthapolous]:
@@
must be satisfied. One can rewrite this as $\sigma(1,\mu)=2S(1,\mu)+2\alpha$ for some constant $\alpha$. That the horizon is locally flat near the axis of symmetry (equivalently, the horizon is topologically $\mathrm{S}^{2}$ [LocalToroidalXanthapolous]) puts a constraint on the odd coefficients
@@
The algebra obtained through the Heisenberg-like boundary condition for local spherical black holes would not differ either from the algebra of spherical black holes considered in [Grumiller-2020]. This is because the topology of the horizon determines the algebra, as mentioned earlier.
@@
Following Chandrasekhar, Xanthopoulos worked out all axially symmetric toroidal local black holes in vacuum in similar coordinates [LocalToroidalXanthapolous]
@@
These equations are different from the spherical-topology case (i.e., equation (34)). In particular, the right-hand side here does not contain any linear derivative of $S$. As before, $\sigma$ is constant on the horizon $\eta=1$, which we will call $2\alpha$ again. The conditions for horizon-regularity are different [LocalToroidalXanthapolous]
@@
We first work out the closed form of $[\xi(T_{1},\Phi_{1}^{A}),\xi(T_{2},\Phi_{2}^{A})]$, i.e., the Lie bracket algebra of the asymptotic Killing vector fields starting from the equation (6). Since now the arbitrary functions $T$ and $\Phi^{A}$ are functions on a flat torus, we can expand them in Fourier modes as follows:
@@
To summarize, the algebra is the semi-direct sum of diffeomorphisms at the space-like section of the horizon (that is, $\mathrm{S}\times\mathrm{S}$ for the toroidal) generated by $\mathcal{J}_{A}$, and an Abelian current generated by $\mathcal{P}^{(s)}$. Note that each of the pairs $\mathcal{P}^{(s)}_{kl}$, $\mathcal{J}^{2}_{mn}$ and $\mathcal{P}^{(s)}_{kl}$, $\mathcal{J}^{3}_{mn}$ forms a BMS3 sub-algebra [Grumiller-2020]. This can easily be verified by mapping
@@ 4 Thermodynamics
Thus, for any nonzero value of $s$ we do not obtain the usual first law [^5] The difference in the prefactors in the two variations in (68) seems to have been overlooked in [Grumiller-2020], which may have contributed to the conclusion that the first law holds.. One can try redefining the Hawking temperature as
@@
compared to equation (13) [^7] This can be written as $Q=\frac{1}{16\pi}\int\mathrm{d}^{d-2}x\left[2T\sqrt{\Omega}+2\Phi^{A}\frac{\sqrt{\Omega}}{{\kappa}}\left(\partial_{t}M_{A}-R_{A}\right)\right]$, and is similar to equation (2.32) in [Donnay:2016], and equation (14) in [Donnay:2019], expect for slight notational differences.
@@
Instead of rescaling the temperature, one can rescale the Hamiltonian factor to preserve the first law. This can be done by redefining the chemical potential $\mathcal{N}^{(s)}$ as $\kappa\left(r+1\right)^{1/(r+1)}\left(\mathcal{P}^{(s)}\right)^{-r/(r+1)}$. However, in this case, the total Hamiltonian will no longer generate unit-time translations. The other problem is that this would introduce the spin parameter in the Hamiltonian (even when written in terms of $\mathcal{P}$), meaning that the $s\neq 0$ soft-hair terms would contribute to the energy. Thus, we find no immediate way to reconcile the first law for nonzero $s$ while maintaining constant surface gravity. It may still be possible to reconcile it with constant $\kappa$ under certain special thermodynamic configurations. For example, as a purely speculative idea, the rescaled temperature could represent the local temperature of an isothermal cavity in a heat bath (see, for example, [York-Brown:1992, Akbar-Gibbons-2003, Akbar:2010]). Whether such an interpretation is possible remains unclear and is left for future investigation.
@@ 5 Conclusion
Near-horizon symmetries can be approached in several ways, subject to different boundary conditions, all giving supertranslations and superrotations but generally differing in their algebras. For isolated holes, the algebra was found to be two copies of Witt and an abelian current [Donnay:2015, Donnay:2016, Donnay:2019]. By formulating a suitable boundary condition in the ADM formalism ingeniously parameterized by a parameter, it was subsequently found that one can obtain the BMS algebra. While the BMS group is necessarily the symmetry group on $\mathscr{I}$ (which is topologically $\mathrm{S}^{2}\times\mathbb{R}$) for asymptotically flat spacetimes, there is no a priori reason for this to be the preferred group for near-horizon symmetries of black holes even when the topology of the horizon is $\mathrm{S}^{2}\times\mathbb{R}$. Thus, this result came as a welcome surprise [^8] As long as the conserved charges on the horizon outnumber the charges at infinity, we have a satisfactory scattering process in the sense of [Hawking:2016]..
All explicit examples previously studied in connection with near-horizon symmetries have strictly (geometrical) spherical horizons, with the only exceptions being the C-metrics [Anabalon2021], which have axially symmetric horizons that are topological two-spheres. The only study of non-spherical topology in near-horizon symmetries to date examines black holes with an $\mathrm{S}^{1}\times\mathrm{S}^{2}$ horizon topology in five dimensions [Giribet2023]. In this paper, we have examined static black holes in four-dimensional vacuum general relativity that are topologically spherical and toroidal but geometrically axially symmetric. None of these solutions, while solving the vacuum Einstein equations, are asymptotically flat. These solutions are regular everywhere, and their horizons are non-singular, unlike the vacuum C-metric, and together they constitute the complete set of axisymmetric static black holes that solve the vacuum Einstein equations. We have found that the symmetry algebra for the toroidal black holes changes to being two copies of the spin-$s$ BMS3 from spin-$s$ BMS4 for spherical topology black holes. For $s=0$, this gives two noncommuting Witt algebras for the toroidal holes.
@@ Appendix A Toroidal Algebra
Here we present detailed calculations for one of the Lie brackets for the toroidal black holes (54). Consider the Lie bracket $[\mathcal{Y}^{2}_{pq},\mathcal{Y}^{3}_{mn}]$. Making use of equation (4), we immediately see $T_{12}=0$. The other two relations work out to be
[…]
@@
$[\mathcal{Y}^{2}_{pq},\mathcal{Y}^{3}_{mn}]=-\mathrm{i}q\mathrm{e}^{\mathrm{i}\left[(p+m)\theta+(q+n)\phi\right]}\partial_{\theta}+\mathrm{i}m\mathrm{e}^{\mathrm{i}\left[(p+m)\theta+(q+n)\phi\right]}\partial_{\phi}=-\mathrm{i}q\mathcal{Y}^{2}_{p+m,q+n}+\mathrm{i}m\mathcal{Y}^{3}_{p+m,q+n}~{}.$ (72)

### v2

@@ 3 Local Axisymmetric Black Holes
Chandrasekhar obtained all vacuum spherical topology axisymmetric black holes in Schwarzschild-like coordinates [30, 31]:
@@
must be satisfied. One can rewrite this as $\sigma(1,\mu)=2S(1,\mu)+2\alpha$ for some constant $\alpha$. That the horizon is locally flat near the axis of symmetry (equivalently, the horizon is topologically $\mathrm{S}^{2}$ [31]) puts a constraint on the odd coefficients
@@
The algebra obtained through the Heisenberg-like boundary condition for local spherical black holes would not differ either from the algebra of spherical black holes considered in [20]. This is because the topology of the horizon determines the algebra, as mentioned earlier.
@@
Following Chandrasekhar, Xanthopoulos worked out all axially symmetric toroidal local black holes in vacuum in similar coordinates [31]
@@
These equations are different from the spherical-topology case (i.e., equation (34)). In particular, the right-hand side here does not contain any linear derivative of $S$. As before, $\sigma$ is constant on the horizon $\eta=1$, which we will call $2\alpha$ again. The conditions for horizon-regularity are different [31]
@@
We first work out the closed form of $\xi(T_{1},\Phi_{1}^{A}),\xi(T_{2},\Phi_{2}^{A})]$, i.e., the Lie bracket algebra of the asymptotic Killing vector fields starting from the equation (6). Since now the arbitrary functions $T$ and $\Phi^{A}$ are functions on a flat torus, we can expand them in Fourier modes as follows:
@@
To summarize, the algebra is the semi-direct sum of diffeomorphisms at the space-like section of the horizon (that is, $\mathrm{S}\times\mathrm{S}$ for the toroidal) generated by $\mathcal{J}_{A}$, and an Abelian current generated by $\mathcal{P}^{(s)}$. Note that each of the pairs $\mathcal{P}^{(s)}_{kl}$, $\mathcal{J}^{2}_{mn}$ and $\mathcal{P}^{(s)}_{kl}$, $\mathcal{J}^{3}_{mn}$ forms a BMS3 sub-algebra [20]. This can easily be verified by mapping
@@ 4 Thermodynamics
Thus, for any nonzero value of $s$ we do not obtain the usual first law [^7] The difference in the prefactors in the two variations in (68) seems to have been overlooked in [20], which may have contributed to the conclusion that the first law holds.. One can try redefining the Hawking temperature as
@@
compared to equation (13) [^9] This can be written as $Q=\frac{1}{16\pi}\int\mathrm{d}^{d-2}x\left[2T\sqrt{\Omega}+2\Phi^{A}\frac{\sqrt{\Omega}}{{\kappa}}\left(\partial_{t}M_{A}-R_{A}\right)\right]$, and is similar to equation (2.32) in [18], and equation (14) in [19], expect for slight notational differences.
@@
Instead of rescaling the temperature, one can rescale the Hamiltonian factor to preserve the first law. This can be done by redefining the chemical potential $\mathcal{N}^{(s)}$ as $\kappa\left(r+1\right)^{1/(r+1)}\left(\mathcal{P}^{(s)}\right)^{-r/(r+1)}$. However, in this case, the total Hamiltonian will no longer generate unit-time translations. The other problem is that this would introduce the spin parameter in the Hamiltonian (even when written in terms of $\mathcal{P}$), meaning that the $s\neq 0$ soft-hair terms would contribute to the energy. Thus, we find no immediate way to reconcile the first law for nonzero $s$ while maintaining constant surface gravity. It may still be possible to reconcile it with constant $\kappa$ under certain special thermodynamic configurations. For example, as a purely speculative idea, the rescaled temperature could represent the local temperature of an isothermal cavity in a heat bath (see, for example, [46, 47, 48]). Whether such an interpretation is possible remains unclear and is left for future investigation.
@@ 5 Conclusion
Near-horizon symmetries can be approached in several ways, subject to different boundary conditions, all giving supertranslations and superrotations but generally differing in their algebras, depending on the spacetime dimension and the topology of the black holes. For isolated holes, the algebra was found to be two copies of Witt and an abelian current [17, 18, 19]. By formulating a suitable boundary condition in the ADM formalism ingeniously parameterized by a parameter, it was subsequently found that one can obtain the BMS algebra. While the BMS group is necessarily the symmetry group on $\mathscr{I}$ (which is topologically $\mathrm{S}^{2}\times\mathbb{R}$) for asymptotically flat spacetimes, there is no a priori reason for this to be the preferred group for near-horizon symmetries of black holes even when the topology of the horizon is $\mathrm{S}^{2}\times\mathbb{R}$. Thus, this result came as a welcome surprise [^10] As long as the conserved charges on the horizon outnumber the charges at infinity, we have a satisfactory scattering process in the sense of [16]..
All explicit examples previously studied in connection with near-horizon symmetries (in four dimensions) have strictly (geometrical) spherical horizons, with the only exceptions being the C-metrics [21], which have axially symmetric horizons that are topological two-spheres. The only study of non-spherical topology in near-horizon symmetries to date examines black holes with an $\mathrm{S}^{1}\times\mathrm{S}^{2}$ horizon topology in five dimensions [22]. In this paper, we have examined static black holes in four-dimensional vacuum general relativity that are topologically spherical and toroidal but geometrically axially symmetric. None of these solutions, while solving the vacuum Einstein equations, are asymptotically flat. These solutions are regular everywhere, and their horizons are non-singular, unlike the vacuum C-metric, and together they constitute the complete set of axisymmetric static black holes that solve the vacuum Einstein equations. We have found that the symmetry algebra for the toroidal black holes changes to being two copies of the spin-$s$ BMS3 from spin-$s$ BMS4 for spherical topology black holes. For $s=0$, this gives two noncommuting Witt algebras for the toroidal holes.
Intuitively, one may expect that algebra, being local, would not be affected by the topology of the horizon. However, as we have seen in this paper, the infinite-dimensional near-horizon symmetry group involves arbitrary functions defined on the horizon, and the space of allowed functions on a torus differs from that on a sphere. This causes the generators to depend on the topology of the horizon, resulting in different algebras. It would be interesting to investigate how topology affects near-horizon symmetries in a more general setting. For example, one could study higher-genus black holes in four (or higher) dimensions in the presence of a negative cosmological constant [49]. This is left for future work.
@@ Appendix A Toroidal Algebra
Here we present detailed calculations for one of the Lie brackets for the toroidal black holes (54). Consider the Lie bracket $\mathcal{Y}^{2}_{pq},\mathcal{Y}^{3}_{mn}]$. Making use of equation (4), we immediately see $T_{12}=0$. The other two relations work out to be
$\Phi^{\theta}_{12}$ $=\Phi^{B}_{1}\partial_{B}\Phi^{\theta}_{2}-\Phi^{B}_{2}\partial_{B}\Phi^{\theta}_{1}=\cancelto{0}{\mathrm{e}^{\mathrm{i}(p\theta+q\phi)}\partial_{\theta}\Phi^{\theta}_{2}}-\mathrm{e}^{\mathrm{i}(m\theta+n\phi)}\partial_{\phi}\mathrm{e}^{\mathrm{i}(p\theta+q\phi)}~,$ (71)
$\Phi^{\phi}_{12}$ $=\Phi^{B}_{1}\partial_{B}\Phi^{\theta}_{1}-\Phi^{B}_{2}\partial_{B}\Phi^{\phi}_{1}=\mathrm{e}^{\mathrm{i}(p\theta+q\phi)}\partial_{\theta}\mathrm{e}^{\mathrm{i}(m\theta+n\phi)}-\cancelto{0}{\mathrm{e}^{\mathrm{i}(m\theta+n\phi)}\partial_{\phi}\Phi^{\theta}_{1}}~.$
@@
$\mathcal{Y}^{2}_{pq},\mathcal{Y}^{3}_{mn}]=-\mathrm{i}q\mathrm{e}^{\mathrm{i}\left[(p+m)\theta+(q+n)\phi\right]}\partial_{\theta}+\mathrm{i}m\mathrm{e}^{\mathrm{i}\left[(p+m)\theta+(q+n)\phi\right]}\partial_{\phi}=-\mathrm{i}q\mathcal{Y}^{2}_{p+m,q+n}+\mathrm{i}m\mathcal{Y}^{3}_{p+m,q+n}~.$ (72)